OpenStudy (anonymous):

find x to three significant digits 2.5^(x/2)=59

OpenStudy (kropot72):

Take logs of both sides: \[\frac{x}{2}\log_{} 2.5=\log_{} 59\] Then divide both sides by log 2.5 \[\frac{x}{2}=\frac{\log_{} 59}{\log_{} 2.5}=you\ can\ calculate\]

OpenStudy (anonymous):

ok thanks

OpenStudy (kropot72):

You're welcome :)

OpenStudy (anonymous):

so x=8.9?

OpenStudy (kropot72):

Good work! Your answer is correct :)

OpenStudy (kropot72):

But remember to state the answer to three significant digits.